### v2

# Accurate close interactions of Stokes spheres using lubrication-adapted image systems

Abstract
Stokes flows with near-touching rigid particles induce near-singular lubrication forces under relative motion, making their accurate numerical treatment challenging. With the aim of controlling the accuracy with a computationally cheap method, we present a new technique that combines the method of fundamental solutions (MFS) with the method of images. For rigid spheres, we propose to represent the flow using Stokeslet proxy sources on interior spheres, augmented by lines of image sources adapted to each near-contact to resolve lubrication. Source strengths are found by a least-squares solve at contact-adapted boundary collocation nodes. We include extensive numerical tests, and validate against reference solutions from a well-resolved boundary integral formulation. With less than 60 additional image sources per particle per contact, we show controlled uniform accuracy to three relative digits in surface velocities, and up to five digits in particle forces and torques, for all separations down to a thousandth of the radius. In the special case of flows around fixed particles, the proxy sphere alone gives controlled accuracy. A one-body preconditioning strategy allows acceleration with the fast multipole method, hence close to linear scaling in the number of particles. This is demonstrated by solving problems of up to 2000 spheres on a workstation using only 700 proxy sources per particle.
Key words: Method of fundamental solutions, method of images, elliptic PDE, Stokes flow, collocation
1 Introduction
The Stokes equations constitute a linearization of the Navier-Stokes equations, valid in the limit of small Reynolds numbers, where viscous forces are dominant over inertial forces. Consequently, the fluid adjusts instantly to any applied force without damping, as the mass of the fluid does not affect the fluid motion. Such fluid systems with immersed particles have a wide range of applications, from the design of a lab-on-a-chip for diagnostics or drug delivery, to water purification and sedimentation processes. In biology, the flow of red blood cells in our veins, the motion of swimming bacteria [1, 2, 3] and the mechanics of the cytoskeleton in our cells, where a complex network of microtubules, actin filaments and molecular motors provide structural support and enable movement within cells [4, 5], can be understood based on a Stokesian model [6, 7]. In materials science, the flow of fibers in e.g. the paper making industry can be characterized [8], and novel materials developed [9, 10].
We focus on the Stokes resistance boundary value problem (BVP) for a suspension of rigid spheres immersed in a given background flow $\boldsymbol{u}_{\infty}$. Let $\Omega\coloneqq\bigcup_{k=1}^{P}\Omega^{(k)}$ denote a collection of $P$ balls, and $\partial\Omega$ denote the union of all of their sphere boundaries $\partial\Omega^{(k)}$. We assume that $\boldsymbol{u}_{\infty}$ is a homogeneous Stokes solution throughout $\mathbb{R}^{3}$, and that in the presence of the spheres, the flow field becomes $\boldsymbol{u}_{\infty}+\boldsymbol{u}$. The unknown disturbance flow field, $\boldsymbol{u}$, therefore solves the BVP
[…]
$\boldsymbol{u}(\boldsymbol{x})$ $\to\boldsymbol{0},$ $\|\boldsymbol{x}\|\to\infty.$
Here, $p$ is the disturbance pressure field, and $\mu$ the constant viscosity of the fluid. We assume no-slip boundary conditions, with given translational and angular velocities for each sphere. For sphere $k$, this means that the velocity boundary data is prescribed as
$\boldsymbol{u}_{\text{bc}}^{(k)}(\boldsymbol{x})=\boldsymbol{v}^{(k)}+\boldsymbol{\omega}^{(k)}\times(\boldsymbol{x}-\boldsymbol{c}^{(k)})-\boldsymbol{u}_{\infty}(\boldsymbol{x}),\quad\boldsymbol{x}\in\partial\Omega^{(k)},$ (2)
with $\boldsymbol{v}^{(k)}$ its given translational velocity, $\boldsymbol{\omega}^{(k)}$ its given angular velocity, and $\boldsymbol{c}^{(k)}$ its center. Aside from solving for the above flow field, it is also of interest to extract the net force and torque on each particle.
As particles move relative to each other in Stokes flow, their interactions become increasingly difficult to resolve accurately with decreasing particle separations, due to growing lubrication forces [11]. Small gaps and strong lubrication come with high resolution requirements and a high computational cost. This is intractable for simulations of systems with many particles or of systems with small Brownian particles, where many realizations are needed to draw statistical conclusions. With simulation methods based on a volumetric grid, such as finite element methods, the narrow gap between particles becomes very costly to mesh. For a boundary integral method (BIE), on the other hand, it is enough to discretize only the particle surfaces, but two other challenges have to carefully addressed: As one solves for a so called layer-density [^1] The simplest interpretation of the layer density is that of a force density over the surface, which holds if a single layer formulation is used in the boundary integral formulation. over the particle surfaces, near-singularities in integrals over the surfaces must be accurately evaluated [12, 13, 14, 15, 16]. Additionally, strong lubrication forces manifest themselves as a sharply peaked layer density in regions close to touching with another particle, which, to be resolved, requires a very fine surface grid, resulting in large linear systems with many unknowns to solve. There is hence a need for more efficient techniques.
Several approximate methods exist to capture lubrication [17, 18]. This is often done by explicitly adding a contribution to particle pair-interactions to compensate for what is not captured with the “regular” computational method or grid [19, 20, 21]. The most famous of these approximate correction strategies is the widely used Stokesian dynamics [22, 23, 24, 25]. The surfaces of the particles are then not explicitly represented and the interaction between particles is governed by a multipole expansion. The corrected hydrodynamic resistances between particles are assumed to be pair-wise additive, which is an approximation that holds only for small particle gaps [26]. Including more terms in the multipole expansion can provide higher accuracy also for moderately separated particles, but implies an increased algebraic complexity. Stokesian dynamics can be accelerated with fast summation techniques both with and without Brownian motion, which enable simulations of large particle systems over long time-spans to a manageable cost [25]. An alternative is the rigid multiblob method [27, 28], where a particle is described as a collection of spherical blobs or beads constrained to move as a rigid body and the interaction between blobs is regularized by the Rotne-Prager-Yamakawa tensor. In [28], we improve on the accuracy of the multiblob method in the far-field and also reduce errors for close interactions. Neither with Stokesian dynamics nor the rigid multiblob method is the accuracy however possible to control. Using a BIE method equipped with special quadrature, controllable accuracy can be obtained for moderately separated particles [13, 14]. One such technique based on spherical harmonics enables fast large-scale simulations for spheres specifically [29, 30]. However, since they are not spatially adaptive, such methods demand an excessive number of surface unknowns at small separation distances, and thus become inefficient.
In the description above, it has been assumed that closely interacting particles also move relative to each other; however, only under relative motion do lubrication forces diverge with vanishing separation distance between particles [11, p. 175]. For bodies at fixed locations subject to a background flow (e.g., a porosity simulation), the singularities are much weaker, so that a non-adaptive BIE method is adequate [13]. The difference in difficulty to resolve problems with or without relative motion has been exploited numerically. In [20], velocities for pairs of particles are split into relative and rigid body motion. Only relative motions then require special treatment as a consequence of a peaked force density over the particle surfaces, while joint rigid body motions are easy to resolve [20]. In [21], lubrication-corrections for particle pairs, similar to those in Stokesian dynamics, are constructed to be frame-invariant, meaning that only relative particle motions are affected by the correction.
When formulating an integral equation, a free space Green’s function is typically used, and boundary conditions are enforced by solving for the layer density in the solution ansatz. If one can instead build a Green’s function that satisfies the boundary conditions, the solution is easily constructed. Such domain specific Green’s functions can sometimes be found using an image or reflection technique, the simplest case being that of a homogeneous Dirichlet condition on a flat plane for the Laplace equation. Such a Laplace image system with known strengths and locations can also be built for two interacting conducting discs or spheres, but here the number of reflections, or image sources, is infinite (this problem was solved 170 years ago by Lord Kelvin [31]; also see [32, §172–§173] and [33, 34]). With more objects, there is an exponential explosion of multiple reflections—even for three disks or spheres the image points form a fractal [35]—rendering a numerical image sum impossible.
Cheng & Greengard [33] considered closely spaced conducting discs (the Laplace equation) in two dimensions, followed by an extension to three dimensions for conducting spheres by Cheng [34], and dielectric spheres by Gan et al. [36]. Here, interactions of the objects get increasingly difficult to resolve with decreasing separation distances: a standard numerical solution in a multipole (Rayleigh expansion) basis centered on each object then demands a large and wasteful number of unknowns per body. The authors introduce a hybrid method which adds to each multipole basis function its (truncated) pairwise image reflection series in each sufficiently close neighbor, thus keeping the number of unknowns per body small regardless of the closeness of separation. Yet, image methods are significantly more complicated for the Stokes case. For a flat wall with a no-slip condition, analytic formulae were first derived for the Stokeslet by Blake [37], and later in a different form by Gimbutas et al. [38]. The methodology for deriving the image system for a Stokeslet in the exterior of one sphere, that enforces a no-slip condition on the surface of the sphere, was introduced in [39, 40], and expressions are found in [41] and [11, §10.2]. They involve multiple types of singularities: the Stokeslet, rotlet, stresslet, and potential dipole, including continuous distributions along the line from image point to center. Although various approximations exist for two distant spheres [11, Ch. 10-11], we do not believe that an exact formula for multiple reflections in two nearby spheres would be practical. Thus a Stokes generalization of the above hybrid method seems out of reach. However, we will sacrifice analytic formulae in favor of a more numerical approach that still draws inspiration from Lord Kelvin, Cheng & Greengard, and the Stokes image for a single sphere.
The approach of this paper is, broadly speaking, a method of fundamental solutions (MFS). The solution flow $\boldsymbol{u}$ is written as a linear combination of fundamental solutions to the Stokes PDE, located in the non-physical region (the sphere interiors), whose coefficients are found by enforcing the boundary data in the least-squares sense at a similar number of surface collocation points. This idea has a long history, and sometimes also goes under the names of the charge simulation method, the method of auxiliary sources, or proxy-sources [42, 43, 44, 45, 46, 47, 48]. A choice of locations and collocation points leading to small errors depends on the domain and data, is heuristic, and is not always possible to determine. However, there are rigorous convergence results in certain smooth geometries, such as the disk with source locations on a concentric circle [49], and when the solution may be analytically continued as a PDE solution up to the source curve [50]. The MFS has the disadvantage of involving ill-conditioned rectangular linear systems, when compared against the well-conditioned square ones arising from discretizing 2nd-kind Fredholm BIEs. In special geometries such as close-to-touching spheres, however, we will see that the advantage of much smaller numbers of unknowns at a given accuracy makes the MFS rather favorable compared to BIEs.
Two sources of numerical errors affect a dynamic simulation of particles in a non-Brownian Stokesian suspension: the error in the Stokes BVP, and the time-step discretization error in evolving the particle coordinates. To properly capture the strong lubrication forces in a dynamic simulation of close-to-touching particles, the time-step size often has to be extremely small, which adds substantially to the computational cost. Larger time steps reduce the computational effort but can lead to particle overlaps. In addition, at very small separation distances, the surface roughness of physical particles becomes non-negligible, and the Stokes equations are no longer a good model for their interactions [51, 52]. To avoid investing a huge computational effort in this regime, a contact-avoiding strategy is needed, such that a minimum particle separation is enforced [7, 53, 30, 54]. In the present work, focus is on solving the Stokes BVP, and time-stepping is left for future work. However, with any future time-stepping strategy in mind, its resolution requirements determine what we here consider a physically relevant particle separation for which we seek accurate solutions.
(a) Surface velocities.
(b) Relative residual at the particle surfaces.
Figure 1: By using an adaptive number of additional source points per closely interacting particle pair, lubrication forces between 20 closely interacting spheres in a cluster (many separated by $10^{-3}$ radii) are resolved with our new technique. Each sphere travels with a random translational and angular velocity. The target accuracy of $10^{-3}$ in the relative residual is reached everywhere on the particle surfaces.
1.1 Contribution and outline
We present an efficient numerical scheme for the Stokes resistance BVP compatible with linear-scaling fast solution algorithms. We target 3-digit (0.1%) uniform error level in the surface velocities, chosen to be small enough that other sources of error discussed above would dominate in practical settings. Moreover, we will show that controllable accuracy can be obtained at a competitive cost for any separation down to $10^{-3}$ radii, far smaller than has been shown in other multi-sphere Stokes solvers [13, 55, 25, 30, 56, 57]. Our cost is higher than that of more approximate techniques such as Stokesian dynamics or the rigid multiblob method, but lower than that of a BIE method accurate to the same level. Our technique is a variant of the MFS discussed above, hence avoids volume discretization, and allows evaluation of potentials right up to the boundary without specialized techniques. To tackle ill-conditioning, we apply so-called one-body preconditioning [46, 50] to create a well-conditioned square system amenable to FMM-accelerated iterative solution in quasi-linear time. This allows large numbers of spheres to be handled on a single workstation. The MFS source locations, types, and collocation nodes must be carefully chosen to capture the near singular force densities involved. One of our main contributions is a recipe for this—inspired by the analytic image distributions discussed above—that adapts to each pairwise particle separation. We propose Stokeslets located on interior concentric “proxy” spheres, complemented by point sources along segments of lines connecting the centers of close-to-touching pairs. We call the latter image points, as their locations discretize an image line source. See Figure 2. The resulting number of unknowns needed to achieve the target accuracy is much less than for a BIE-method, as we show in our experiments. The result is a higher efficiency scheme.
A small-scale example of the type of problem we solve is visualized in Figure 1 with a cluster of 20 closely interacting unit spheres with relative motion. It is a challenging problem: each sphere is $\delta=10^{-3}$ from at least one neighbor, with velocities set to randomly sampled rigid body motion, as visualized in panel 1(a). With the number of source points per particle set adaptively based on separation distances, panel 1(b) shows that the target accuracy of $10^{-3}$ is reached everywhere on the particle surfaces.
Remark 1 (Recent Stokes MFS work).
Since the preprint of this paper appeared on arXiv, another MFS Stokes paper appeared by Jordan & Lockerby [58]. They present results for the resistance and mobility problems, including nonspherical particles. However, in contrast to our work, only problems with weak lubrication effects (large separations) are treated. They apply Gauss–Seidel iterations to account for particle-particle effects, giving a $\mathcal{O}(P^{2})$ cost, contrasting the linear scaling obtained via the FMM in our work. Results in [58] are shown for particles along a line, with interactions greatly simplified beyond a certain distance. In contrast, we present a method with full interactions and test it with space-filling configurations.
The outline of the paper is as follows. The Stokes MFS is introduced in Section 2, including the basic discretization, the one-body preconditioning scheme, and the enhanced MFS image sources needed for problems with strong lubrication forces. The parameter selection for the basic MFS discretization is discussed in Section 3, and the resulting performance for large-scale problems with small lubrication effects is discussed in Section 4. For the more challenging problems involving strong lubrication forces, fundamental solutions at the additional source points have to be chosen such that their combined approximation power is sufficient for resolving the solution to the set of PDEs. The types of fundamental solutions needed at the image points are investigated in Section 5.1, together with distributions of additional collocation points to capture the near-singular lubrication flow. The investigations lead to our adaptive scheme for the number of image points, depending on the particle-particle separations only, which is then applied to problems with clusters of closely interacting spheres affected by lubrication in Section 5.2. To demonstrate that a target accuracy of $10^{-3}$ can be reached for all physically relevant particle separations, errors in the flow fields, forces and torques computed with MFS are assessed by comparing to reference solutions computed with a BIE-method equipped with special quadrature, by self-convergence tests, and by measuring the residual at the particle surfaces at points different from the collocation points. Our findings are summarized in Section 6.
2 The method of fundamental solutions
In this section, we first introduce the basic discretization that will be used in the paper, followed by a description of the preconditioning strategy that we employ. In subsection 2.3, we discuss how to enhance this basic discretization to allow for accurate solutions also in the case of strong lubrication forces, i.e. for particle interactions with relative motion at small separation distances.
2.1 The basic discretization for one particle
Possibly the simplest solution to the PDE and decay condition at infinity in (1) is given by
$\boldsymbol{u}(\boldsymbol{x})=\mathbb{S}(\boldsymbol{x}-\boldsymbol{y})\boldsymbol{f},$ (3)
valid in $\mathbb{R}^{3}\backslash\{\boldsymbol{y}\}$. The tensor $\mathbb{S}$ is the Stokes fundamental solution known as the Stokeslet, and $\boldsymbol{f}$ is a point force located at $\boldsymbol{y}$. In free-space, without confinements, defining $\boldsymbol{r}\coloneqq\boldsymbol{x}-\boldsymbol{y}$, $\boldsymbol{I}\in\mathbb{R}^{3\times 3}$ the identity matrix, and $\left(\boldsymbol{r}\boldsymbol{r}^{T}\right)_{ij}=r_{i}r_{j}\in\mathbb{R}^{3\times 3}$ the dyadic product, the Stokeslet is
$\mathbb{S}(\boldsymbol{r})=\frac{1}{8\pi\mu\|\boldsymbol{r}\|}\left(\boldsymbol{I}+\frac{\boldsymbol{r}\boldsymbol{r}^{T}}{\|\boldsymbol{r}\|^{2}}\right).$ (4)
There is of course a pressure solution $p$ accompanying $\boldsymbol{u}$ in (3), but we will not need it. To represent the boundary condition $\boldsymbol{u}_{\text{bc}}$ on a single particle, however, a linear combination of fundamental solutions at $N$ source points $\boldsymbol{y}_{j}$ in the particle interior is needed, so that
$\boldsymbol{u}(\boldsymbol{x})=\sum_{j=1}^{N}\mathbb{S}(\boldsymbol{x}-\boldsymbol{y}_{j})\boldsymbol{\lambda}_{j}.$ (5)
Without loss of generality we fix a sphere radius $R=1$. One imposes boundary conditions via collocation at a set of points $\boldsymbol{x}_{i}$, $i=1,\dots,M$, covering the surface, where $M>N$. This gives the overdetermined linear system
$\sum_{j=1}^{N}\mathbb{S}(\boldsymbol{x}_{i}-\boldsymbol{y}_{j})\boldsymbol{\lambda}_{j}\;=\;\boldsymbol{u}_{\text{bc}}(\boldsymbol{x}_{i}),\qquad i=1,\dots,M,$ (6)
which is solved in the least-squares sense to find the set of strength vectors $\boldsymbol{\lambda}_{j}$. In the basic version of our algorithm, the $N$ source points $\boldsymbol{y}_{j}$ cover a concentric interior proxy-sphere of radius $R_{p}<1$. As source and collocation points live on separated surfaces, the singularities in the Stokeslet at $\boldsymbol{x}=\boldsymbol{y}_{j}$, and thus the singular quadratures needed for BIE methods, are avoided. The tall rectangular matrix $\boldsymbol{G}$ of size $3M\times 3N$, whose $3\times 3$ block elements are $\boldsymbol{G}_{ij}\coloneqq\mathbb{S}(\boldsymbol{x}_{i}-\boldsymbol{y}_{j})$, we refer to as the target-from-source matrix.
In order to minimize the boundary condition residual and the ill-conditioning, we seek as uniform as possible a point distribution on the physical sphere and proxy sphere, respectively. The many available quasi-uniform point distributions include Fibonacci grids [59, 60], HEALpix grids [61], quadratures with a (possibly different) equispaced grid on each line of latitude, and Gaussian quadrature schemes for spherical harmonics such as Lebedev points (based on the octahedron) or Ahrens–Beylkin points (based on the icosahedron); see [62] and references within. However, we recommend so-called “spherical design points” [63] for surface- and proxy-points, since with constant weights they integrate spherical harmonics up to a certain degree $p$ exactly, even though their efficiency (they need $N\approx p^{2}/2$ points) is not quite as high as Gaussian ($N\approx p^{2}/3$). The constancy of weights means that they are quasi-uniform, unlike variable-weight quadrature schemes which are more clustered; we find that this improves stability. [^2] Note that certain of the above grids could enable exploiting azimuthal symmetry, but for the basic proxy discretization the SVD costs are too small for it to matter. Spherical designs only exist for certain $N$ values. We use points computed [^3] These are available in double-precision accuracy for $N$ up to around 16000 at https://web.maths.unsw.edu.au/~rsw/Sphere/EffSphDes/. by Womersley [64]. In our tests, proxy-surfaces of radii $R_{p}$ are each discretized with $N$ spherical design nodes, while collocation surfaces are discretized with $M\approx 1.2N$ spherical design nodes, as in Figure 2(a).
The ratio $M/N\approx 1.2$ was found by a convergence study, and is typical for MFS applications [44, 50]. We do not know of rigorous analysis of this ratio, apart from the case $M/N=1$ for the 2D Laplace equation [49]. Intuitively, a ratio greater than $1$ is beneficial because the potential generated by somewhat nearby proxy points contains frequency components higher than merely the Nyquist frequency associated with their spacing.
Remark 2 (Conditioning).
The condition number of the target-from-source matrix for a single sphere is expected to grow exponentially like $(1/R_{p})^{c\sqrt{N}}$ for some constant $c$ close to unity. To see this heuristically, note that [29, (3.9)–(3.11)] show that the continuous single-layer operator from a source sphere of radius $R_{p}<1$ to a concentric sphere of radius 1 has each eigenfunction a linear combination of vector spherical harmonics within a given degree $n$ and order $m$, with eigenvalue $R_{p}^{n}$ (up to algebraic prefactors). Furthermore, $N$ spherical design points accurately discretize all harmonics of degree $n\leq p\approx c\sqrt{N}$, for some constant $c$ close to 1. Thus the minimum eigenvalue of the matrix $\boldsymbol{G}$ is expected to be $R_{p}^{c\sqrt{N}}$, and the maximum eigenvalue $\mathcal{O}(1)$, justifying the claim. Numerically, the decay of singular values matches this model, for $c\approx 1$, as visualized in Figure 18(a) (Appendix A.2). With a typical choice $R_{p}=0.7$, choosing $N$ in the range $500$ to $2000$ gives $\kappa(\boldsymbol{G})$ in the range $10^{5}$ to $10^{9}$.
The effects of the parameters $N$ (proxy-points per sphere) and $R_{p}$ (proxy-radius) are investigated in Section 3. We also give numerical examples that show that the ill-conditioning of the MFS system matrix does not pose a problem, as long as the norm of the resulting solution vector $\boldsymbol{\lambda}$ is bounded, and that the latter is insured by setting $R_{p}$ large enough to enclose any singularities (i.e. images) in the continuation of the solution into the interior of the sphere. This is standard for the MFS in other settings [44, 50].
[…]
(a) Basic discretization
Refer to caption Line of images $\beta_{2}$$\beta_{1}$$\delta$ Proxy- surface
(b) With image-enhancement: sketch
[…]
(c) With image-enhancement: point distributions
Figure 2: Panel (a) displays the basic discretization, with the original source points $\boldsymbol{y}_{j}$, $j=1,\dots,N$, on the radius $R_{p}$ proxy-surface in blue, and collocation points $\boldsymbol{x}_{i}$, $i=1,\dots,M$, in black on the unit sphere surface. Both source and collocation points are obtained from spherical design nodes. Panels (b) and (c) display the image-enhanced grid of source and collocation points for a pair of close-to-touching spheres. Additional image points are distributed along a line segment in the interior of each sphere, while additional collocation points are distributed “above” each line of images. The collocation points are located on two spherical caps determined by angles $\beta_{1}$ and $\beta_{2}$ surrounding the point of closest approach to another particle, as in panel (b). If a particle is closely interacting with more than one neighbor, one region of extra source and collocation points of this type is added for every near-contact. In panel (c), collocation points are indicated in black and image points in red, generated via (26)-(27).
2.2 One-body preconditioned iterative solution for many particles
Next, we describe a technique for solving the large overdetermined and ill-conditioned system for the unknown MFS coefficients $\boldsymbol{\lambda}$ on $P$ particles. Let $\boldsymbol{x}^{(k)}_{i}$, $i=1,\dots,M$, denote the collocation points on particle $k$, and $\boldsymbol{y}^{(k)}_{j}$, $j=1,\dots,N$ denote the sources in its interior. Denoting by
$\boldsymbol{X}=\begin{bmatrix}{\boldsymbol{x}^{(1)}}^{T}&{\boldsymbol{x}^{(2)}}^{T}&\dots&{\boldsymbol{x}^{(P)}}^{T}\end{bmatrix}^{T}$
the stacked vector of collocation points for all particles, we seek to solve the overdetermined linear system
$\boldsymbol{G}\boldsymbol{\lambda}=\boldsymbol{b},\qquad\text{with }\quad\boldsymbol{b}=\boldsymbol{u}_{\text{bc}}(\boldsymbol{X}),$ (7)
in the least-squares sense. Here, the entire target-from-source matrix $\boldsymbol{G}\in\mathbb{R}^{3MP\times 3NP}$ has blocks $\boldsymbol{G}^{(kk^{\prime})}\in\mathbb{R}^{3M\times 3N}$, each representing the effect on the surface velocities $\boldsymbol{u}$ of body $k$ due to the Stokeslet sources within body $k^{\prime}$, giving the block form
$\boldsymbol{G}=\begin{bmatrix}\boldsymbol{G}^{(11)}&\boldsymbol{G}^{(12)}&\dots&\boldsymbol{G}^{(1P)}\\ \boldsymbol{G}^{(21)}&\boldsymbol{G}^{(22)}&\ldots&\ldots\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{G}^{(P1)}&\ldots&\ldots&\boldsymbol{G}^{(PP)}\end{bmatrix}.$ (8)
A dense direct least squares solution would be unfeasibly slow for large $P$, with a cost $\mathcal{O}\left(MP\cdot(NP)^{2}\right)$. However, the ill-conditioning inherited from that of the MFS for each particle prevents an iterative solution in the form (7). To address this we use one-body right preconditioning. The diagonal block $\boldsymbol{G}^{(kk)}$ takes the form of the matrix in (6), with sources at $\boldsymbol{y}^{(k)}_{j}$, $j=1,\dots,N$, in body $k$ evaluated at the collocation points $\boldsymbol{x}^{(k)}_{i}$, $i=1,\dots,M$, on its surface. For identical particles considered here, denote this one-body block by $\boldsymbol{B}$. The block-diagonal preconditioning matrix contains the blocks $\boldsymbol{B}^{+}$, and hence represents the pseudo-inverses of the target-from-source matrices for each body in isolation. This turns the ill-conditioned rectangular system with $3N$ unknowns per body into a well-conditioned square system with $3M$ unknowns per body, which is solved via GMRES. This idea was used by Stein & Barnett in [50], and first considered in an MFS Helmholtz problem by Liu & Barnett in [46]. It has not to our knowledge been proposed in the 3D Stokes MFS setting before.
One-body preconditioning has subtleties, so we now present it in detail. For each particle $k$, we work with the unknown coefficient vector $\boldsymbol{\mu}^{(k)}:=\boldsymbol{B}\boldsymbol{\lambda}^{(k)}\in\mathbb{R}^{3M}$, representing surface collocation velocities, instead of its source strength vector $\boldsymbol{\lambda}^{(k)}\in\mathbb{R}^{3N}$. Substituting $\boldsymbol{\lambda}^{(k)}=\boldsymbol{B}^{+}\boldsymbol{\mu}^{(k)}$ directly into (7) would result in diagonal blocks $\boldsymbol{B}\boldsymbol{B}^{+}\in\mathbb{R}^{3M\times 3M}$ that are rank deficient because $M>N$, rather than the identity matrix $\boldsymbol{I}\in\mathbb{R}^{3M\times 3M}$ that expresses the true surface velocity self-interaction. Thus we replace each such diagonal block by $\boldsymbol{I}$, and recover a well-conditioned formulation. The preconditioned linear system to solve, with square system matrix $\hat{\boldsymbol{G}}$ of size $\mathbb{R}^{3MP}\times\mathbb{R}^{3MP}$, is then
$\hat{\boldsymbol{G}}{\boldsymbol{\mu}}:=\begin{bmatrix}\boldsymbol{I}&\boldsymbol{G}^{(12)}\boldsymbol{B}^{+}&\dots&\boldsymbol{G}^{(1P)}\boldsymbol{B}^{+}\\ \boldsymbol{G}^{(21)}\boldsymbol{B}^{+}&\boldsymbol{I}&\ldots&\ldots\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{G}^{(P1)}\boldsymbol{B}^{+}&\ldots&\ldots&\boldsymbol{I}\end{bmatrix}\begin{bmatrix}\boldsymbol{\mu}^{(1)}\\ \boldsymbol{\mu}^{(2)}\\ \vdots\\ \boldsymbol{\mu}^{(P)}\end{bmatrix}={\begin{bmatrix}\boldsymbol{b}^{(1)}\\ \boldsymbol{b}^{(2)}\\ \vdots\\ \boldsymbol{b}^{(P)}\end{bmatrix}.}$ (9)
Taking the SVD, $\boldsymbol{B}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}$, and the pseudo-inverse $\boldsymbol{B}^{+}$ is
$\boldsymbol{B}^{+}=\boldsymbol{V}\boldsymbol{\Sigma}^{+}\boldsymbol{U}^{T},$ (10)
where the matrix $\boldsymbol{\Sigma}^{+}$ is computed as a pseudo-inverse in the sense that singular values below a threshold $\epsilon_{\text{trunc}}$ are truncated, so that the elements of $\boldsymbol{\Sigma}^{+}$ are limited in magnitude. However, the matrix $\boldsymbol{B}^{+}$ cannot be formed and then applied in a backward stable fashion [65]; two factors must be stored and applied sequentially to prevent catastrophic cancellations. The matrix in (9) could therefore not be built explicitly, even if speed and storage limitations were ignored. Instead, to perform the matrix-vector multiply with $\hat{\boldsymbol{G}}$ in the GMRES iteration to solve (9), we propose the following three steps:
  1. 1.
    Apply the self-interaction pseudo-inverse $\boldsymbol{B}^{+}$ backward-stably (using a pair of dense matrix-vector multiplies as in [65, 66, 50, 67]) to the current coefficient vector for each particle,
    $\hat{\boldsymbol{\lambda}}^{(k)}=\boldsymbol{V}\boldsymbol{\Sigma}^{+}\left({\boldsymbol{U}}^{T}\boldsymbol{\mu}^{(k)}\right),\quad k=1,\dots,P.$ (11)
  2. 2.
    Compute, either using direct summation or the fast multipole method (FMM), the all-to-all interaction
    $\boldsymbol{u}_{\text{All}}=\boldsymbol{G}\hat{\boldsymbol{\lambda}},$ (12)
    between sources and collocation points for all particles, recalling (8).
  3. 3.
    Subtract off the incorrect self-interactions and replace them by the effect of identity blocks, to give
    $\hat{\boldsymbol{G}}{\boldsymbol{\mu}}=\boldsymbol{u}_{\text{All}}-\boldsymbol{B}_{\text{blkdiag}}\hat{\boldsymbol{\lambda}}+\boldsymbol{\mu},$ (13)
    where $\boldsymbol{B}_{\text{blkdiag}}$ is the block-diagonal matrix with diagonal blocks $\boldsymbol{B}$.
If direct summation is used, of course steps 2 and 3 could be combined, but in order to tackle the largest problems we will focus on the use of the FMM where the steps must be separate. Upon GMRES convergence, the vector $\boldsymbol{\mu}$ is again via (11) transformed back to the source strength vector $\boldsymbol{\lambda}$. The latter can then be used for solution velocity or pressure field evaluation throughout the fluid.
2.3 Image enhancement for problems with strong lubrication forces
For challenging problems with strong lubrication, Stokeslets on interior proxy-spheres cannot provide enough resolution unless $R_{p}$ is made very close to $1$ and the number of proxy-sources $N$ is made impractically large. It is much more efficient to keep $N$ and $R_{p}$ small while adding other types of fundamental solutions at locations informed by image methods discussed in the introduction. Hence, only with the basic discretization does $N$ denote the total number of vector valued sources.
Due to the linearity of the set of PDEs in (1), not only the Stokeslet in (4), but also differentiated Green’s functions are fundamental solutions that satisfy the set of equations everywhere except at the source point. The most basic ones are the rotlet, $\mathbb{R}$, stresslet, $\mathbb{T}$ and potential dipole, $\mathbb{D}$, defined in the following. We start by the gradient of the Stokeslet, the force dipole tensor [68, 69]. The symmetric part of this tensor can be written in terms of the rank-3 tensor known as the stresslet, $\mathbb{T}$, with [69]
$\mathbb{T}_{ijk}(\boldsymbol{r})=-\frac{1}{8\pi\mu}\frac{3r_{i}r_{j}r_{k}}{\|\boldsymbol{r}\|^{5}}$ (14)
and a flow field represented by the stresslet can be expressed as
$u_{i}(\boldsymbol{x})=\mathbb{T}_{ijk}(\boldsymbol{x}-\boldsymbol{y})C_{jk}.$ (15)
Here, Einstein’s summation convention is used and $\boldsymbol{C}$ is a constant strength matrix for the stresslet located at $\boldsymbol{y}$ [69]. The anti-symmetric part of the force dipole tensor is the rotlet, $\mathbb{R}$, with
$\mathbb{R}(\boldsymbol{r})=\frac{1}{8\pi\mu}\frac{1}{\|\boldsymbol{r}\|^{3}}\begin{bmatrix}0&r_{3}&-r_{2}\\ -r_{3}&0&r_{1}\\ r_{2}&-r_{1}&0\end{bmatrix}.$ (16)
The associated flow field is given by
$\boldsymbol{u}(\boldsymbol{x})=\mathbb{R}(\boldsymbol{x}-\boldsymbol{y})\boldsymbol{t},$ (17)
where $\boldsymbol{t}$ is a point torque at $\boldsymbol{y}$. The fourth type of singularity, the potential dipole, $\mathbb{D}$, can be interpreted as an opposite source and sink of fluid mass separated by an infinitesimal distance along a given unit vector [69]. A potential dipole with strength $\boldsymbol{d}$ at $\boldsymbol{y}$ gives rise to a fluid velocity field of the form
$\boldsymbol{u}(\boldsymbol{x})=\mathbb{D}(\boldsymbol{x}-\boldsymbol{y})\boldsymbol{d},$ (18)
where
$\mathbb{D}(\boldsymbol{r})=\frac{1}{4\pi}\left(-\frac{\boldsymbol{I}}{\|\boldsymbol{r}\|^{3}}+3\frac{\boldsymbol{r}\boldsymbol{r}^{T}}{\|\boldsymbol{r}\|^{5}}\right).$ (19)
With $\mathbb{S}$, $\mathbb{T}$, $\mathbb{R}$ and $\mathbb{D}$ in place, a generalization of the representation in (5) is to write the solution to (1) due to a single particle as
$u_{i}(\boldsymbol{x})=\sum_{\alpha=1}^{N_{\text{S}}}\mathbb{S}_{ij}(\boldsymbol{x}-\boldsymbol{y}_{\alpha}^{\text{S}})f_{\alpha,j}+\sum_{\alpha=1}^{N_{\text{T}}}\mathbb{T}_{ijk}(\boldsymbol{x}-\boldsymbol{y}_{\alpha}^{\text{T}})C_{\alpha,{jk}}+$ (20)
$+\sum_{\alpha=1}^{N_{\text{R}}}\mathbb{R}_{ij}(\boldsymbol{x}-\boldsymbol{y}_{\alpha}^{\text{R}})t_{\alpha,j}+\sum_{\alpha=1}^{N_{\text{D}}}\mathbb{D}_{ij}(\boldsymbol{x}-\boldsymbol{y}_{\alpha}^{\text{D}})d_{\alpha,j},$
where the singularity locations, $\{\boldsymbol{y}_{\alpha}^{\text{S}}\}_{\alpha=1}^{N_{\text{S}}}$, $\{\boldsymbol{y}_{\alpha}^{\text{T}}\}_{\alpha=1}^{N_{\text{T}}}$, $\{\boldsymbol{y}_{\alpha}^{\text{R}}\}_{\alpha=1}^{N_{\text{R}}}$ and $\{\boldsymbol{y}_{\alpha}^{\text{D}}\}_{\alpha=1}^{N_{\text{D}}}$, may vary with singularity type.
Remark 3 (Stresslet strengths).
All the terms in the sums in (20) are of the form of a matrix-vector product, except for the terms involving the stresslet, $\mathbb{T}_{ijk}(\boldsymbol{x}-\boldsymbol{y}_{\alpha}^{\text{T}})C_{\alpha,{jk}}$. This can however be accomplished by restricting the unknown strength matrix $\boldsymbol{C}_{\alpha}$ to be of the form $C_{\alpha,{lm}}=h_{\alpha,l}q_{\alpha,m}$, for some known vector $\boldsymbol{q}_{\alpha}\in\mathbb{R}^{3}$ and unknown $\boldsymbol{h}_{\alpha}\in\mathbb{R}^{3}$.
The numbers $N_{\text{S}},N_{\text{T}},N_{\text{R}}$ and $N_{\text{D}}$ may vary per particle, since particles have different numbers of neighbors, each at different distances. To indicate this, we use $N^{(k)}$ to denote the total number of vector source unknowns for particle $k$, with
$N^{(k)}=N^{(k)}_{\text{S}}+N^{(k)}_{\text{T}}+N^{(k)}_{\text{R}}+N^{(k)}_{\text{D}}.$ (21)
As with our basic discretization, strengths for the different source types are determined in the least-squares sense via collocation at the particle surface (we postpone a precise description of collocation point parameters to Section 5.1). This results in each matrix block $\boldsymbol{G}^{(kk^{\prime})}$ in (8) being replaced by a horizontal stack of four matrix blocks, one for each type of unknown in (20), to give
$\boldsymbol{G}^{(kk^{\prime})}=\bigl{[}\boldsymbol{G}_{\text{S}}^{(kk^{\prime})},\boldsymbol{G}_{\text{T}}^{(kk^{\prime})},\boldsymbol{G}_{\text{R}}^{(kk^{\prime})},\boldsymbol{G}_{\text{D}}^{(kk^{\prime})}\bigr{]}.$
Explicitly, the elements in each block are $\boldsymbol{G}_{\text{S},ij}^{(kk^{\prime})}:=\mathbb{S}(\boldsymbol{x}_{i}^{(k)},\boldsymbol{y}_{j}^{(k^{\prime})})$, and analogously for types $\mathbb{R}$ and $\mathbb{D}$, while stresslets instead have elements $\boldsymbol{G}_{\text{T},ij}^{(kk^{\prime})}:=\mathbb{T}(\boldsymbol{x}_{i}^{(k)},\boldsymbol{y}_{j}^{(k^{\prime})})\cdot\boldsymbol{h}^{(k^{\prime})}_{j}$. The full bare system matrix $\boldsymbol{G}$ is now of size $3\sum_{k=1}^{P}M^{(k)}$ by $3\sum_{k=1}^{P}N^{(k)}$, and the diagonal self-interaction blocks $\boldsymbol{G}^{(kk)}$ in the preconditioning strategy of Section 2.2 (previously all denoted by $\boldsymbol{B}$) are now in general different.
We now turn to the placement of these different types of fundamental solutions along the center-center line for each pair of close-to-touching particles. Our strategy is motivated by the reflection of points in spheres in the Laplace case, which we now review. The reflection of a point $\boldsymbol{x}$ exterior to a unit sphere with center $\boldsymbol{c}^{(k)}$ gives the image point [31, 34]
$\boldsymbol{x}^{\prime}=\boldsymbol{c}^{(k)}+\frac{\left(\boldsymbol{x}-\boldsymbol{c}^{(k)}\right)}{\|\boldsymbol{x}-\boldsymbol{c}^{(k)}\|^{2}}.$ (22)
The (unique) decaying exterior Laplace BVP solution for boundary data due to a charge at $\boldsymbol{x}$ is given by the potential from a (modified) charge at $\boldsymbol{x}^{\prime}$. In other words, the continuation of this BVP solution into the sphere, as a Laplace solution, has a singularity only at $\boldsymbol{x}^{\prime}$. This has two consequences for the MFS for exterior BVPs (e.g. [49]): i) such a singularity (if close to the surface) controls the convergence rate, and ii) if all such singularities are not enclosed (“shielded”) by the proxy surface, the MFS coefficient norm will grow exponentially, causing a numerical round-off instability [70, 44].
Consider now two disjoint unit spheres centered at $\boldsymbol{c}^{(1)},\boldsymbol{c}^{(2)}$ and separated by a small distance $\delta$, as illustrated in Figure 3. Starting with the center coordinates of each of the spheres reflected in the other sphere, and then computing reflections of reflections by applying (22) iteratively, an infinite sequence of image points is formed in each sphere, lying on a single line as indicated in the figure. The image points accumulate at $\boldsymbol{x}_{\text{acc}}^{(1)}$ in sphere 1 and at $\boldsymbol{x}_{\text{acc}}^{(2)}$ in sphere 2. For a given $\delta$, these points may be solved for as fixed points of the iteration, giving
$\boldsymbol{x}_{\text{acc}}^{(1)}$ $=\frac{\boldsymbol{c}^{(1)}+\boldsymbol{c}^{(2)}}{2}-\sqrt{\delta+\delta^{2}/4}\cdot\frac{\boldsymbol{c}^{(2)}-\boldsymbol{c}^{(1)}}{\|\boldsymbol{c}^{(2)}-\boldsymbol{c}^{(1)}\|},$ (23)
$\boldsymbol{x}_{\text{acc}}^{(2)}$ $=\frac{\boldsymbol{c}^{(1)}+\boldsymbol{c}^{(2)}}{2}+\sqrt{\delta+\delta^{2}/4}\cdot\frac{\boldsymbol{c}^{(2)}-\boldsymbol{c}^{(1)}}{\|\boldsymbol{c}^{(2)}-\boldsymbol{c}^{(1)}\|}.$
As $\delta\to 0$, these accumulation points lie a distance ${\mathcal{O}}(\delta^{1/2})$ inside the surface, which would ruin the convergence and stability of the MFS when using concentric proxy spheres. Attempting to augment the MFS with the true multiple-reflection series (as in Cheng–Greengard [33, 34]) would result in too many images at small $\delta$, and—as described in the introduction—the continuous line-densities generated in the Stokes case render such analytic multiple reflections intractable. Noticing that such a 2-sphere Stokes solution would involve sources only on the lines from centers to the accumulation points motivates our proposal to discretize these lines with a small number of quadrature nodes of unknown strengths.
[…]
Figure 3: Lines of image points obtained by repeated reflections of each sphere center in the other sphere via (22). Accumulation points for the lines of images depend on the separation distance $\delta$ between the two spheres.
For unit spheres discretized with concentric interior proxy-spheres of radius $R_{p}$, we first need a criterion for when additional sources along such an image line out to the accumulation point are needed at all. A simple proxy-sphere of radius $R_{p}$ is stable for MFS (in the sense of having small coefficient norm) only when it contains the accumulation point (and thus all other images). Let
$R_{\text{acc}}(\delta)=1+\delta/2-\sqrt{\delta+\delta^{2}/4}$ (24)
denote the distance of the accumulation point to the center of the sphere. Then, given $R_{p}$, let $\delta^{*}$ denote the critical particle-particle separation where the image accumulation point lies on the sphere with radius $R_{p}$. Precisely, $\delta^{*}$ solves
$R_{\text{acc}}(\delta^{*})=R_{p},$
and it is easy to verify that the unique solution to the above is
$\delta^{*}=\frac{(1-R_{p})^{2}}{R_{p}}.$ (25)
If a pairwise distance $\delta$ is less than $\delta^{*}$, then we will need to place additional image sources which discretize the part of the line from the center to $\boldsymbol{x}_{\text{acc}}$ lying outside of the proxy-sphere. The line segment of interest in the interior of the first particle is hence parameterized by
$\boldsymbol{l}^{(1)}(t)=\boldsymbol{c}^{(1)}+t\frac{\left(\boldsymbol{c}^{(2)}-\boldsymbol{c}^{(1)}\right)}{\|\boldsymbol{c}^{(2)}-\boldsymbol{c}^{(1)}\|},\quad t\in(R_{p},R_{\text{acc}}(\delta)],$ (26)
and can be discretized with $n_{\text{im}}$ nodes $\boldsymbol{l}_{j}^{(1)}=\boldsymbol{l}^{(1)}(t_{j})$, $j=1,\dots,n_{\text{im}}$, with $t_{j}$ given by
$t_{j}=R_{p}(1+\Delta)+\left(R_{\text{acc}}(\delta)-R_{p}(1+\Delta)\right)\cos\left(\frac{(j-1)}{2n_{\text{im}}}\pi\right).$ (27)
The nodes are constructed from half of a set of type-2 Chebyshev nodes, clustered towards (and including) the accumulation point; see Figure 2(c). This mimics a quadrature scheme for the interval. For conditioning reasons, the $n_{\text{im}}$th node is here given an offset from the proxy-surface via a small distance $\Delta$, which we fixed at $\Delta=0.05$.
As an ansatz, a linear combination of $\mathbb{S}$, $\mathbb{T}$, $\mathbb{R}$ and $\mathbb{D}$ will be centered at each such image node, with an efficient combination chosen numerically in Section 5.1. For all particle-particle distances larger than $\delta^{*}$, the Stokeslet sources on the proxy-surface alone will allow an MFS representation with small coefficient norm. Considering the target accuracy of $10^{-3}$, we will in Section 3 choose an $R_{p}$ such that $\delta^{*}<0.3$, meaning that image points are only required for separations less than this.
In addition, the fundamental solutions $\mathbb{T}$, $\mathbb{R}$ and $\mathbb{D}$ are all singular at $\boldsymbol{x}=\boldsymbol{y}_{\alpha}$, similarly as the Stokeslet $\mathbb{S}$. In a neighborhood of the additional clustered source points for every near contact of a sphere, collocation points are therefore refined on the sphere surface, as in Figure 2. This has to be done locally, as any additional collocation points will increase the number of unknowns to solve for in the preconditioned system. In our adaptive scheme presented in Section 5.1, both source and collocation points will be set adaptively based on the particle-particle separation for every neighboring pair.
Remark 4 (Extracting net forces and torques).
Having solved for the source strengths $\boldsymbol{\lambda}$ using preconditioned GMRES, the net force $\boldsymbol{F}^{(k)}$ and torque $\boldsymbol{T}^{(k)}$ on particle $k$ are extracted as follows. Given point forces $\boldsymbol{f}_{\alpha}$ and torques $\boldsymbol{t}_{\alpha}$ corresponding to Stokeslets and rotlets in the particle interior,
$\boldsymbol{F}^{(k)}=\sum_{\alpha=1}^{N_{\text{S}}}\boldsymbol{f}_{\alpha},\quad\boldsymbol{T}^{(k)}=\sum_{\alpha=1}^{N_{\text{R}}}\boldsymbol{t}_{\alpha}+\sum_{\alpha=1}^{N_{\text{S}}}\left(\boldsymbol{y}_{\alpha}^{\text{S}}-\boldsymbol{c}^{(k)}\right)\times\boldsymbol{f}_{\alpha}.$ (28)
Conveniently, this avoids any need to integrate stresses over the sphere surfaces.
3 Parameters for the basic concentric proxy sphere
This section serves to optimize the MFS proxy sphere parameters (the number of proxy-sources $N$ and the radius of the proxy-surface $R_{p}$) used in later experiments. No image points are used, and the number of collocation points is fixed at $M\approx 1.2N$, as discussed in Section 2.
As a quantification of the error, we will here and later use the pointwise relative residual, as defined by
$\epsilon_{\text{res}}(\boldsymbol{x})\coloneqq\|\boldsymbol{u}(\boldsymbol{x})-\boldsymbol{u}_{\text{bc}}(\boldsymbol{x})\|_{2}/\|\boldsymbol{u}_{\text{bc}}(\boldsymbol{x})\|_{2},$ (29)
and
$\epsilon_{\text{res}}^{\max}\coloneqq\max\limits_{\boldsymbol{x}\in\partial\Omega}\left(\epsilon_{\text{res}}(\boldsymbol{x})\right),$ (30)
its maximum approximated using a dense set of points sampled on the particle surfaces. We also sometimes use the maximum force/torque error relative to a reference, as defined by
[…]
with $\boldsymbol{\mathcal{F}}$ a stacked vector of forces and torques for all particles in the system.
(a) Translation or rotation BC: $\epsilon_{\text{res}}^{\max}$.
(b) Translation BC: $\epsilon_{\text{res}}^{\max}$ and $\epsilon_{FT}$, with axis linear in $\sqrt{N}$.
Figure 4: Convergence test for two unit spheres $\delta$ apart, discretized with $N$ source points at inner proxy-surfaces of radius $R_{p}=0.6$, and affected by two types of boundary conditions, translation or rotation, as described in main text. Panel (a) displays rate dependence on $\delta$. In panel (b), dashed curves show convergence as predicted by $\mathcal{O}(R_{\text{acc}}^{\sqrt{N}})$ for the max relative residual $\epsilon_{\text{res}}^{\max}$, while the max force/torque error $\epsilon_{FT}$, relative to a reference solution determined with $N=3529$, converges at a rate of $\mathcal{O}(R_{\text{acc}}^{1.5\sqrt{N}})$.
Throughout this section, the geometrical setup is two unit spheres separated by a distance $\delta$. We start with a convergence test in $N$ with the two particles either rotating relative to each other, with velocities $\boldsymbol{v}^{(1)}=\boldsymbol{0}$, $\boldsymbol{v}^{(2)}=\boldsymbol{0}$, $\boldsymbol{\omega}^{(1)}=\boldsymbol{e}_{z}$ and $\boldsymbol{\omega}^{(2)}=-\boldsymbol{e}_{z}$, or translating together, with $\boldsymbol{v}^{(1)}=\boldsymbol{v}^{(2)}=\boldsymbol{e}_{x}$ and $\boldsymbol{\omega}^{(1)}=\boldsymbol{\omega}^{(2)}=\boldsymbol{0}$. The resulting $\epsilon_{\text{res}}^{\max}$ as function of $N$ is visualized for a few different choices of $\delta$ in Figure 4(a). The convergence appears to be root exponential, i.e. spectral in $p\sim\sqrt{N}$. More specifically, we notice the rate is consistent with $\log(1/R_{\text{acc}})$, where $R_{\text{acc}}$ is given by (24). By analogy with rigorous 2D Laplace MFS theory [49], the rate can be well predicted by the decay rate of spherical harmonic coefficients induced by a Stokes singularity at the radius $R_{\text{acc}}$ from the previous section. One also must assume that the maximum harmonic degree captured by the MFS is $p\approx\sqrt{N}$, corresponding to one proxy point per spherical harmonic term (as in Remark 2). However, the convergence of $\epsilon_{FT}$ is a bit faster: [^4] We leave for future investigations to understand the exact convergence rates. Figure 4(b) fits a root-exponential rate of $1.5\log(1/R_{\text{acc}})$. In this test, $R_{p}=0.6$. For larger $R_{p}$ relative to $R_{\text{acc}}$, $R_{p}$ instead starts to control the convergence rate (not shown), due to “aliasing” error in the MFS dominating [44, Thm. 3 and Rmk. 4].
Next, we describe how to optimize the parameter $R_{p}$ in a test where both $R_{p}$ and $\delta$ are varied. In the panels of Figure 5, the critical separation $\delta^{*}$, recalling (25), is drawn in red, as function of $R_{p}$. For a fixed $R_{p}$, $n_{\text{im}}$ image points will later be added when needed to resolve lubrication forces. Looking for a small $\delta^{*}$ goes hand-in-hand with the wish for $R_{p}$ to be as large as possible, for the proxy-surface to better resolve challenging boundary data, and for the MFS coefficients $\boldsymbol{\lambda}$ to be reasonably small in magnitude. Two test cases are considered in this test to motivate that the discretization we choose will make a good choice for $\delta>\delta^{*}$ and general no-slip boundary data:
  1. 1.
    The particles travel with equal and opposite velocities perpendicular to their line of centers (shearing motion). Lubrication forces are known to scale as $\mathcal{O}(\log(1/\delta))$ [11].
  2. 2.
    The particles travel towards each other with equal and opposite velocities, $\boldsymbol{v}^{(1)}=-\boldsymbol{v}^{(2)}$, along their line of centers (squeezing motion). Lubrication forces are known to scale as $\mathcal{O}(1/\delta)$ [11], and the problem is hence more challenging to resolve than case 1. The resulting net force on each of the particles can be compared to an analytical expansion derived by Brenner, where [71]
    $\|\boldsymbol{F}^{(k)}\|=8\pi\|\boldsymbol{v}^{(k)}\|\sinh\alpha\sum_{n=1}^{\infty}\frac{n(n+1)}{(2n-1)(2n+3)}\times\left[\frac{4\cosh^{2}(n+1/2)\alpha+(2n+1)^{2}\sinh^{2}\alpha}{2\sinh(2n+1)\alpha-(2n+1)\sinh 2\alpha}-1\right],$ (32)
    and $\alpha$ is defined by $\cosh\alpha=1+\delta/2$ with $\alpha\sim\sqrt{\delta}$ as $\delta\to 0$.
To proceed, we fix the number of points at the proxy-surface to $N=686$ or $N=1353$ (the largest number of precomputed spherical design nodes smaller than $N=700$ and $N=1400$ respectively). The ratio of the smallest to largest singular values of the self-interaction target-from-source matrix are all above $\epsilon_{\text{mach}}$ and no truncation is therefore in use in the one-body preconditioning scheme of Section 2.2. The resulting $\epsilon_{\text{res}}^{\max}$, as given by (30), are visualized in Figures 5(a), 5(b) and 5(e) as functions of $R_{p}$ and $\delta$. Shearing motion in panel 5(a) is easier to resolve than squeezing motion in 5(b) and 5(e), and a difference in error levels can also be noted between the two grid resolutions, represented in panels 5(b) and 5(e) for the same test problem. This is as predicted by MFS theory: with a larger number of source points, aliasing errors, dominating for larger $R_{p}$, are decreased. Figure 5(d) visualizes $\|\boldsymbol{\lambda}\|_{\infty}$ determined with the finer grid with $N=1353$. From panels 5(a), 5(b), 5(d) and 5(e), note that a large $R_{p}$ corresponds to a small $\delta^{*}$, while a smaller $R_{p}$ generally gives a smaller relative residual $\epsilon_{\text{res}}^{\max}$ for a fixed particle separation $\delta$, however at the expense of a larger coefficient vector (as predicted by MFS theory). Note however that, due to round-off, such big coefficient norms typically cause loss of accuracy. The error in the computed force on one of the particles is shown in Figures 5(c) and 5(f). Note that the relative error in the force is generally smaller than the maximum relative residual, often by a few orders of magnitude.
[…]
(a) $N=686$, shearing motion, $\epsilon_{\text{res}}^{\max}$.
[…]
(b) $N=686$, squeezing motion, $\epsilon_{\text{res}}^{\max}$.
[…]
(c) $N=686$, squeezing motion, $\epsilon_{\text{FT}}$.
[…]
(d) $N=1353$, squeezing motion, maximum coefficient magnitude, $\|\boldsymbol{\lambda}\|_{\infty}$.
[…]
(e) $N=1353$, squeezing motion, $\epsilon_{\text{res}}^{\max}$.
[…]
(f) $N=1353$, squeezing motion, $\epsilon_{\text{FT}}$.
Figure 5: A pair of unit spheres discretized with $N$ sources on proxy-surfaces of radius $R_{p}$ are separated by a small distance $\delta$. A sweep over $R_{p}$ and $\delta$ is made to investigate accuracy and MFS coefficient magnitudes. The test is done for two settings: case 1 (squeezing motion), and case 2 (shearing motion). Squeezing motion is harder to resolve, confirmed by comparing panels (a) and (b). The error in the computed force relative to the results of Brenner in (32), denoted by $\epsilon_{\text{FT}}$, is a more forgiving measure of the error than the maximum relative residual $\epsilon_{\text{res}}^{\max}$; compare panels (c) and (b) for $N=686$ and (f) and (e) for $N=1353$. For a given $R_{p}$, additional image points can only be added for $\delta<\delta^{*}$ (to the left of the red curve representing $\delta^{*}(R_{p})$ in each panel). For large $\delta$, $\epsilon_{\text{FT}}$ and $\epsilon_{\text{res}}^{\max}$ are capped at $10^{-6}$ which is the tolerance chosen for GMRES. Vertical black lines are drawn for reference at $\delta=5\cdot 10^{-2}$, $\delta=1\cdot 10^{-1}$, $\delta=5\cdot 10^{-1}$ and $\delta=1$.
For the coarser grid and given a general target accuracy of $10^{-3}$, we pick $R_{p}=0.63$, for which $\delta^{*}\approx 0.271$. With $N=1353$, $R_{p}$ can be chosen larger for the same target accuracy. We set $R_{p}=0.7$, where $\delta^{*}\approx 0.129$. The coarse grid is used throughout the numerical experiments in this paper, while the dense grid will serve as a reference in some of the tests. By extrapolating from the data in Figure 5, it can be concluded that to resolve the interaction for particles with relative motion and a small separation distance $\delta\ll 0.1$, the number of sources on the proxy-surfaces would have to be intractably large. For such interactions, we will instead add image points as given by (26)-(27).
4 Large scale examples with moderate lubrication forces
In this section we show the performance of the basic MFS scheme using uniform proxy spheres and no lubrication-adapted images. This can reach the target 3-digit accuracy if the particles are well-separated (Example 1 below), or have no relative motion (Example 2), or both.
We consider two example geometries. With no images, every particle is discretized equally, which means that only one SVD is needed to compute the particle pseudo-inverse used repeatedly to build the preconditioned linear system in (9). To evaluate $\boldsymbol{u}_{\text{All}}$ in (12), and to evaluate the fluid flow in the exterior domain in a post-processing step, linear scaling is achieved using the MATLAB interface to the Stokes kernels in the multithreaded FMM3D library [72, 73]. In both examples in this section, the GMRES tolerance is set to $10^{-7}$ and the FMM tolerance is set to $10^{-8}$. The implementations are in MATLAB R2022b, using an 8-core workstation with Intel® Xeon® CPU E5-2620 @ 2.00GHz and 128 GB of RAM.
Example 1.
First, consider a dilute suspension of spheres placed randomly in a layer, with geometry as visualized in Figure 6(b). With the discretization choice $N=686$ and $R_{p}=0.63$ from Section 3, the well-separated condition means $\delta>\delta^{*}\approx 0.271$, recalling (25). To generate the suspension geometry, the position of each new sphere is drawn from a uniform distribution in a box of dimension $[L,L,2]$ until the distance to any neighbor is at least $\delta=\delta^{*}$. The boundary conditions are generated by random angular and translational velocities for every particle. The computed flow magnitude in a plane cutting through an example suspension containing 2000 spheres is displayed in Figure 6(b). A maximum relative residual $\epsilon_{\text{res}}^{\max}$, recalling (30), matching the target tolerance of $10^{-3}$ is obtained, as visualized in Figure 6(d). To determine $\epsilon_{\text{res}}^{\max}$, the residual is evaluated in 500 points per surface, different from the set of collocation points used to enforce boundary conditions. The force/torque error $\epsilon_{\text{FT}}$, determined relative to a reference computed with $N=1353$ via (31), is nearly three digits more accurate than $\epsilon_{\text{res}}^{\max}$. The cost of FMM is linear in the number of particles $P$, as illustrated in Figure 6(d), but the number of GMRES iterations grows weakly, which is the reason why the reported total solution time in Figure 6(d) is slightly superlinear. The solution time is roughly 10 s per particle. For comparison, the time to compute the single particle pseudo-inverse with an SVD is 2.5 s.
(a) Suspension geometry with 2000 spheres.
[…]
(b) Flow magnitude in a plane cutting through the configuration of spheres in (a). White space is occupied by particles.
(c) The error $\epsilon_{FT}$ in the force/torque vector determined with $N=686$ sources per particle is computed relative to a fine grid reference determined with $N=1353$. Displayed are also $\epsilon_{\text{res}}^{\max}$ for both choices of $N$.
(d) The computational cost per GMRES iteration scales linearly with the number of particles, but the iteration count grows weakly, giving a slightly worse than linear total solution time.
Figure 6: Example with a dilute suspension of spheres in a layer.
Example 2.
Next, consider a cluster of fixed spheres in a background flow, as exemplified in Figure 7(a) for a shear flow and in Figure 7(b) for a uniform flow. The cluster is grown so that each sphere is a distance of $\delta$ away from at least one other sphere, where $\delta=10^{-2}$ or $\delta=10^{-3}$. The spheres are fixed in space and do not move with the fluid, meaning that the total velocity field at the sphere surfaces is zero. The boundary condition for the disturbance field due to the presence of the particles in (2) in shear flow can be expressed as
$\boldsymbol{u}_{\text{bc}}(\boldsymbol{x})=-\left(\boldsymbol{E}_{\infty}\boldsymbol{x}+\boldsymbol{\omega}_{\infty}\times\boldsymbol{x}\right),\,\quad\text{with }\quad\boldsymbol{E}_{\infty}=0.5\begin{bmatrix}0&\dot{\gamma}&0\\ \dot{\gamma}&0&0\\ 0&0&0\end{bmatrix},\,\quad\boldsymbol{\omega}_{\infty}=\begin{bmatrix}0\\ 0\\ -0.5\dot{\gamma}\end{bmatrix},$ (33)
and $\dot{\gamma}=5$ the shear rate. For the uniform flow, we have
$\boldsymbol{u}_{\text{bc}}(\boldsymbol{x})=-\boldsymbol{u}_{\infty},\quad\text{ with }\quad\boldsymbol{u}_{\infty}=\begin{bmatrix}5&0&0\end{bmatrix}^{T}.$ (34)
An $\epsilon_{\text{res}}^{\max}$ smaller than the target accuracy of $10^{-3}$ can be obtained independently of the choice of $\delta$ or the number of particles for both flows, presented in Table 1 for the shear case.
Table 1: Number of GMRES iterations to convergence, maximum relative residuals $\epsilon_{\text{res}}^{\max}$ and maximum coefficient magnitudes $\|\boldsymbol{\lambda}\|_{\infty}$ for simulations of $P$ particles in a shear flow (Figure 7(a)), using $N=686$. The GMRES tolerance is set to $10^{-6}$ and the FMM tolerance is $10^{-8}$.
Setting Quantity
GMRES iters. $\boldsymbol{\epsilon}_{\text{res}}^{\max}$ $\|\boldsymbol{\lambda}\|_{\infty}$
$\boldsymbol{P=100}$
$\delta=10^{-2}$, $R_{p}=0.75$ 92 $1.1\cdot 10^{-4}$ $9.83$
$\delta=10^{-3}$, $R_{p}=0.75$ 122 $9.6\cdot 10^{-5}$ $9.52$
$\delta=10^{-3}$, $R_{p}=0.65$ 116 $6.9\cdot 10^{-5}$ $316$
$\boldsymbol{P=2000}$
$\delta=10^{-2}$, $R_{p}=0.75$ 144 $9.3\times 10^{-5}$ $18.3$
$\delta=10^{-3}$, $R_{p}=0.75$ 176 $1.3\times 10^{-4}$ $18.9$
$\delta=10^{-3}$, $R_{p}=0.65$ 156 $1.5\times 10^{-4}$ $1013$
[…]
(a) Shear flow
[…]
(b) Uniform flow
Figure 7: Streamlines around clusters of 2000 fixed spheres in a background flow. Each sphere is separated by $\delta=10^{-3}$ from at least one neighbor and discretized with $N=686$ proxy-sources, resulting in a total number of $4.8$ million degrees of freedom in the preconditioned linear system. No image points are needed as the particles do not move relative to each other. The solutions have about 4-digit accuracy (uniform boundary condition error), and take around 10 hours on one workstation.
The results for the uniform flow are similar, but with slightly faster convergence. The FMM dominates the cost with about 90% of the total solution time, and is more dominant for larger systems. For the largest suspensions with 2000 particles, the preconditioned systems solved via FMM have $4.81\times 10^{6}$ degrees of freedom. In comparison, applying the pseudo-inverse and computing the self-interaction (steps 1 and 3 in the algorithm described in Section 2.2) each takes about 5% of the solution time. The total CPU times for $P=2000$ are 12.8 h and 10.1 h for shear and uniform flow, while $P=100$ runs in 0.51 h and 0.45 h, respectively. Note in Table 1 that the smallest particle-particle distance, $\delta=10^{-3}$, is, surprisingly, no harder to resolve than the larger particle-particle distance $\delta=10^{-2}$. This is a consequence of the fact that the force density—the coefficient vector $\boldsymbol{\lambda}$—is benign, despite the closeness of the particles; see the reported $\|\boldsymbol{\lambda}\|_{\infty}$ in the table. This is in contrast to the prior test in Figure 5, where two particles travel relative to each other and $\|\boldsymbol{\lambda}\|_{\infty}$ gets very large for small $\delta$ and any relevant $R_{p}$. This contrast matches the observations in [20], who attribute the difficulty of resolving a peaked force density to problems with relative motion between particles. af Klinteberg & Tornberg [13] in their study of 131 ellipsoids using QBX also implicitly make use of the smoothness of the force density for static particles.
Despite this apparent lack of singularities at $R_{acc}$, we still find that decreasing $R_{p}$ results in a larger coefficient norm, as shown by comparing lines 2–3 or 5–6 of Table 1. The required number of GMRES iterations for clusters of two different sizes are also compared in this table. Similarly as in Example 1, and for different choices of $R_{p}$ and $\delta$, the number of iterations grows only weakly with $P$.
Streamlines illustrated in Figure 7 are computed by propagating a set of tracer particles in the fluid, using Matlab’s ode45 to solve the ODE
$\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{d}t}=\boldsymbol{G}(\boldsymbol{x}_{t})\boldsymbol{\lambda}$ (35)
for the tracer particle locations $\boldsymbol{x}_{t}$ at time $t$. The MFS coefficient vector $\boldsymbol{\lambda}$ is computed once for the fixed particle geometry.
5 Closely interacting particles under relative motion
We now turn to Stokes sphere problems with strong lubrication forces, arising for small separations with relative motions. We show that they are efficiently solved by augmenting the coarse proxy sphere with fundamental solutions placed along discretized image lines defined by (26)-(27). We start by studying which singularity types are needed along these lines, together with a strategy for choosing additional collocation points. The result is an adaptive strategy to set the number of image and collocation points for each pair-wise separation $\delta$. This is demonstrated on clusters of spheres in Section 5.2. The coarse grid with $N=686$ and $R_{p}=0.63$ (from Section 3) is used throughout for the proxy spheres. Finally, in Section 5.3, we discuss the GMRES iteration count for small $\delta$.
One-body right preconditioning (as in Section 2.2) can still be applied with the extra images, but with the difference that the number of vector-valued sources (and collocation nodes) now can vary for each particle, depending on the proximity to its neighbors. The self-interaction blocks of the target-from-source matrix will therefore be different for each particle, increasing the SVD costs. The clustered image and collocation points also cause more ill-conditioning of the self-interaction blocks, as their columns are nearly linearly dependent. This means that truncation in the pseudo-inverses is needed: we set $\epsilon_{\text{trunc}}=5\cdot 10^{-12}$ throughout, via a tradeoff between residual error vs coefficient norm (see Section A in the Appendix). Moreover, since it slightly improves the conditioning of the target-from-source matrix, left-preconditioning (row-scaling) has also been used in the paper whenever image points are utilized, discussed in Remark 5. The decay of the singular values with preconditioning is exemplified in Figure 18(c) in the Appendix, to be compared to the singular values without preconditioning shown in Figure 18(b).
5.1 Choice of fundamental solutions and extra collocation points
We start by numerically testing different combinations of the four types of fundamental solutions at the approximate image points for a single near-contact with separation $\delta$: Stokeslets $\mathbb{S}$, rotlets $\mathbb{R}$, stresslets $\mathbb{T}$ and potential dipoles $\mathbb{D}$. Figure 8(a) compares the convergence with respect to the number of vector unknowns per image line, for a few such combinations, and shows that Stokeslets, rotlets and potential dipoles is the most efficient combination (shown in red). Here, a fixed large number of collocation points, $M=5000$, was used, sufficient to resolve lubrication forces. Figures 8(c) and 8(c) show that adding the converged image line to the coarse proxy sources reduces the residual by seven digits. From now on, the number of image sources per particle per contact will be $3n_{\text{im}}$, accounting for three types of fundamental solutions at the $n_{\text{im}}$ image points $\boldsymbol{l}_{j}$ in (26)-(27).
[…]
(b) No images
(c) With images
[…]
Figure 8: Combinations of fundamental solutions along image lines reduce the maximum relative residual $\epsilon_{\text{res}}^{\max}$ on the particle surface. Panel (a) shows convergence for various combinations, for a pair of spheres separated by a distance of $\delta=10^{-2}$. Panels (b) and (c) plot the relative residual, as defined in (29), over one of the particles, without and with the best combination of fundamental solutions at $n_{\text{im}}=20$ image points: the improvement is dramatic.
In order to minimize the number of collocation nodes $M$ while still resolving the solution, we add spherical caps of uniformly sampled collocation points “above” the line of images, as shown back in Figure 2(b-c). For each cap, two parameters have to be set: the number of points and the angle of the cap, together determining an average node density. For robustness in configurations with more than two close-to-touching particles, we have found that the best results are obtained by superposing two such caps per particle per contact, one of which has a smaller opening angle $\beta_{1}$, and the other a wider angle $\beta_{2}$. The numbers of collocation points added by the two caps on one particle involved in a single near-contact are
$M^{\text{cap}}_{1}=\alpha_{1}\cdot 3n_{\text{im}}\qquad\text{ and }\quad M^{\text{cap}}_{2}=\alpha_{2}\cdot 3n_{\text{im}},$ (36)
with $\alpha_{1}$ and $\alpha_{2}$ the respective upsampling parameters to be chosen. For cap $k=1$ or $2$, we choose a quasi-uniform Fibonacci grid [60] in the spherical coordinates $(\theta_{j},\phi_{j})$ relative to the inter-particle axis, namely
$\theta_{j}$ $=\arccos\left(1-j\cos{\beta_{k}}/M_{k}^{\text{cap}}\right),$ (37)
$\phi_{j}$ $=2j\pi\Phi^{-1}\mod{2\pi},\quad j=1,\dots,M_{k}^{\text{cap}},$
with $\Phi=(1+\sqrt{5})/2$ the golden ratio. The node density has to be sufficiently high in a sufficiently large area above the accumulation point to resolve the boundary data. As before, we target a uniform surface residual of $10^{-3}$. The effect of different choices of $\beta_{1}$ and $\beta_{2}$ is visualized in Figure 9 for a configuration of three spheres forming an equilateral triangle. Here $\delta=10^{-3}$, being the most challenging interaction that we would like to resolve, and we set $n_{\text{im}}=30$. Based on experiments of this type, and to keep the parameter selection simple, we fix for later use $\alpha_{1}=\alpha_{2}=6$, $\beta_{1}=\pi/5$ and $\beta_{2}=\pi/60$, independently of $\delta$.
(a) $\beta_{2}=\pi/200$
(b) $\beta_{2}=\pi/60$
(c) $\beta_{2}=\pi/10$
[…]
Figure 9: The dependence on the density of extra collocation points for the relative surface residual. Collocation points are added on two spherical caps per close contact (see Figure 2) and three different choices of $\beta_{2}$ in (37) are investigated, with $\beta_{1}$ fixed at $\pi/5$. The sphere has two close neighbors ($n_{c}=2$) at a distance $\delta=10^{-3}$ in an equilateral triangle. The borders of the two additional spherical caps are shown with black lines. Here, $M^{\text{cap}}_{1}=M^{\text{cap}}_{2}$ are fixed following (36), with $\alpha_{1}=\alpha_{2}=6$. The density of collocation points has to be sufficiently high in the region above the image lines to resolve the boundary condition. This is the case in panel (b), where the relative residual $\epsilon_{\text{res}}(\boldsymbol{x})$ defined in (29) is below the target accuracy of $10^{-3}$, whereas in panel (a), $\beta_{2}$ is chosen too narrow and in panel (c), $\beta_{2}$ is chosen too wide. In the example, $N=686$, $R_{p}=0.6$, $n_{\text{im}}=30$ and $M=1.2\cdot N+2\cdot 6\cdot n_{c}\cdot 3n_{\text{im}}$.
Remark 5 (Left-preconditioning).
To improve on the conditioning with clustered source and collocation points, the following preconditioning is applied from the left. Motivated by a desire to preserve the $L_{2}$-norm of the solution [74], let $\boldsymbol{W}$ be a diagonal matrix formed by the local area elements corresponding to each collocation point, with the matrix blocks
$\left[\sqrt{w_{1}}\boldsymbol{I}_{3\times 3},\sqrt{w_{2}}\boldsymbol{I}_{3\times 3},\dots,\sqrt{w_{i}}\boldsymbol{I}_{3\times 3},\dots,\sqrt{w_{MP}}\boldsymbol{I}_{3\times 3}\right]$
on the diagonal and $w_{i}$ the local area associated with collocation point $i$. For a particle involved in one near-contact, $w_{i}$ takes one of three values, depending on if the collocation point $\boldsymbol{x}_{i}$ belongs to the basic grid or lies on one of the spherical caps of additional uniformly distributed collocation nodes. The preconditioned problem takes the form
$\boldsymbol{W}\boldsymbol{G}\boldsymbol{\lambda}=\boldsymbol{W}\boldsymbol{b},$ (38)
using the notation from (7), to be solved for $\boldsymbol{\lambda}$ in the least-squares sense. The effect of the preconditioning is that matrix rows corresponding to the two spherical caps of collocation points will be weighted, with their number density taken into account.
We now show that the above parameter choices achieve the target residual error for various boundary conditions. The convergence with $n_{\text{im}}$ is measured for three distinct settings with lubrication:
  1. 1.
    Equal and opposite translational velocities, with the spheres travelling towards each other (squeezing motion). Similarly to the test with squeezing motion without images in Figure 5 in Section 3, the computed forces on the particles are compared to the analytical expansion by Brenner in (32). The maximum relative residuals $\epsilon_{\text{res}}^{\max}$ over the particle surfaces are shown in Figure 10(a), together with the convergence of the relative force errors in Figure 10(d).
  2. 2.
    Equal and opposite translational velocities, with the spheres moving perpendicular to their line of centers (shearing motion), with residual convergence in Figure 10(b).
  3. 3.
    Equal and opposite angular velocities. The spheres rotate around the $z$-axis and are placed next to each other on the $x$-axis, with residual convergence in Figure 10(c).
As convergence occurs in these tests, the coefficient norm drops from roughly $10^{8}$ down to $10^{3}$, possibly interpreted as an insufficient $n_{\text{im}}$ struggling to capture the true singularity line density. The squeezing motion is the most difficult boundary condition to resolve, requiring the largest number of image points to reach the target accuracy of $10^{-3}$. We investigate this further in Figure 11, where a sweep over $n_{\text{im}}$ is made for two spheres undergoing squeezing motion, a variable distance $\delta$ apart. Based on this, we may estimate the number of image points needed to reach a target accuracy of $10^{-3}$, given $N=686$ and $R_{p}=0.63$, as
$n_{\text{im}}(\delta)=\left\{\begin{aligned} \lceil-8.72\log_{10}(\delta)-6.15\rceil,\quad&\delta<0.15,\\ 0,\quad&\text{otherwise.}\end{aligned}\right.$ (39)
This is our adaptive rule for $n_{\text{im}}$, and it will be validated in Section 5.2. For all $\delta>10^{-3}$, note specifically that 20 image points per particle per contact are enough to resolve lubrication to the desired level of accuracy, corresponding to only 60 additional unknown vector source strengths.
Finally, we may apply the above pair-wise recipe to the multi-particle case using the general singularity notation of Section 2.3. Let the $k$th particle have $n_{c}\in\{0,1,\dots,\}$ near contacts, with separations $\delta_{j}$, $j=1,\dots,n_{c}$. Its number of Stokeslets is then $N_{\text{S}}^{(k)}=N+\sum_{j}^{n_{c}}n_{\text{im}}(\delta_{j})$, where the first term accounts for the coarse proxy points. We also have $N_{\text{R}}^{(k)}=N_{\text{D}}^{(k)}=\sum_{j}^{n_{c}}n_{\text{im}}(\delta_{j})$, while $N_{\text{T}}^{(k)}=0$, giving the total number of vector unknowns $N^{(k)}$ as in (21). The associated total number of collocation nodes for particle $k$ is
$M^{(k)}=1.2\cdot N+2\cdot 6\cdot\sum_{j=1}^{n_{c}}3n_{\text{im}}(\delta_{j}),$ (40)
the first term being the coarse nodes and the rest (recalling $\alpha_{1}=\alpha_{2}=6$) the two added node caps.
(a) Squeezing motion.
(b) Shearing motion.
(c) Equal and opposite rotation.
(d) Squeezing motion: force error relative to the analytic result of Brenner in (32).
Figure 10: Two spheres separated by $\delta$ require a different number of image points $n_{\text{im}}$ per particle depending on the boundary conditions; squeezing is the hardest to resolve. In all tests, $n_{\text{im}}=20$ is enough to reach an error level of $10^{-3}$ or less, displayed for the max relative residual $\epsilon_{\text{res}}^{\max}$ in panels (a), (b) and (c), and the force error for squeezing in panel (d).
Figure 11: The number of image points per particle, $n_{\text{im}}$, varied for two spheres a varying distance $\delta$ apart undergoing squeezing motion, discretized with $N=686$ and $R_{p}=0.63$. The red line is the fitted estimate (39) of the number of image points required per particle per close interaction, as a function of $\delta$, to reach the target accuracy $10^{-3}$.
5.2 Validating the pairwise-adaptive strategy in clusters of spheres
We now present four examples with clusters of spheres tightly packed together, where lubrication is resolved by choosing the number of image and collocation points adaptively, following the pairwise image line discretization (39). We will find that simple unions of these pairwise image lines perform just as accurately for multi-sphere clusters, a remarkably useful result (whose explanation may be related to Cheng–Greengard’s three-disk analysis [33, Lem. 3.3]).
Recall that the coarse proxy sphere always has $N=686$ and $R_{p}=0.63$. In self-convergence tests, we compare against results determined with an image discretization tuned for a larger number of sources, where $N=1353$, $R_{p}=0.7$, and 30 image points per particle per near-contact are used. In one of the examples, a BIE reference solution is used to validate the self-convergence error, but also to highlight the difficulties with BIE that we can overcome with the proposed MFS technique. This BIE solution is computed with a double layer formulation equipped with quadrature by expansion (QBX), developed by af Klinteberg & Tornberg [13] and Bagge & Tornberg [14], where the quadrature tolerance is to $10^{-6}$.
Example 3 (A self-convergence test for the exterior flow field).
In the first example, a triangle of spheres a distance $\delta=10^{-3}$ apart move with translational velocities of unit magnitude towards the center of the triangle. The flow field is evaluated in a large set of points in the plane cutting through the spheres, and in a self-convergence test, the maximum relative velocity difference in the exterior domain is $7.6\cdot 10^{-4}$, i.e., below the target accuracy of $10^{-3}$. The flow field is displayed in Figure 12(a), while the relative difference in the flow field is displayed in Figure 12(b).
[…]
Figure 12: Self-convergence test for an example with three unit spheres forming a triangle of side length $2+10^{-3}$. No-slip boundary conditions are applied, with each sphere assigned a translational velocity pointing towards the center of the triangle, resulting in lubrication forces between the particles. In panel (a), the flow field is visualized in the plane of the triangle, with a zoomed-in plot between two of the particles to the right, where the flow velocity magnitude is the largest. In panel (b), the pointwise relative difference in the computed flow field using a coarser and a finer MFS grid is displayed. With the 20 image points per particle per near-contact used in the example, the maximum relative difference is $7.6\cdot 10^{-4}$.
Example 4 (Accuracy gain with adaptive images).
We display the difference in error level with the adaptively set image points compared to the basic MFS with proxy-sources only. The geometry is a tetrahedron of unit spheres, as displayed in Figure 13(a), with varying side length of the tetrahedron, given by $2+\delta$, and the relative force/torque error $\epsilon_{\text{FT}}$, as defined by (31), is quantified for each $\delta$ against the fine MFS grid in Figure 13(b). Errors $\epsilon_{\text{FT}}$ are less than the target tolerance of $10^{-3}$ with the use of images, while without image points, accuracy suffers when $\delta<\delta^{*}$, as expected. For each $\delta$, 20 different problems are solved, with randomly sampled no-slip boundary conditions.
(a) Test geometry in Examples 4 and 5.
(b) Self-convergence test for the geometry in (a).
Figure 13: In panel (a), four unit spheres form a tetrahedron of side length $2+\delta$. Source points on proxy-surfaces are visualized in blue, image points in red and collocation points in black. The picture corresponds to the coarse MFS grid, with $N=686$ and $n_{\text{im}}=20$ image points per particle per near-contact. Extra collocation points are added close to the points of near-contact, following (40). Panel (b) displays a self-convergence test with varying $\delta$. In blue, accuracy is controlled with image and collocation points set based on the estimate (39). In black, without extra source and collocation points, the relative error in the computed force and torque, $\epsilon_{\text{FT}}$ defined by (31), diverges with decreasing $\delta$. For $\delta>\delta^{*}$, no image points are used in either case. Shaded regions display max and min for 20 different randomized boundary conditions.
Example 5 (Comparison of accuracy and efficiency relative to a BIE scheme).
With test geometry as in Example 4, the following example serves two purposes: to validate the self-convergence error against results determined with an established BIE solver equipped with QBX quadrature, and to highlight the ability of the MFS scheme to resolve the interaction for problems where a BIE struggles to reach accuracy  [^5] Since the mobility problem—where forces and torques on the particles are provided then rigid body motions of the particles are computed—generally converges faster than the resistance problem under study, we choose to generate reference solutions with QBX via the mobility formulation; forces and torques are assigned and the computed velocities are used as input to our MFS scheme.. For moderately small $\delta$, where the peaks in the BIE layer density are not as severe, comparisons are possible, while for very small $\delta$, the force density peak of width $\mathcal{O}(\sqrt{\delta})$ [17] requires an impractically fine grid to be resolved in the BIE setting. In addition, the precomputations needed for employing QBX are more costly the finer the resolution. We test two grid-resolutions for use with QBX to demonstrate these claims: a coarser grid with $30\times 30$ discretization points per particle, and a finer grid with $60\times 90$ grid points (5400 vector unknowns). Due to CPU time constraints, the fine QBX-based reference is only determined for two sets of boundary conditions at each $\delta$: a unit torque applied in the $z$-direction on every particle, or a unit force in the $x$-direction on every particle. In Figure 14, the error relative to the fine-grid BIE remains small for smaller values of $\delta$ compared to the coarser BIE-grid, yet both errors start diverging for sufficiently small $\delta$. This onset varies with boundary condition, but above this, comparisons against both the coarse and fine BIE-grids show similar errors as the self-convergence test (which displays the same data as in Example 4), validating the MFS against the BIE solution. For large $\delta>0.3$, the BIE errors match the QBX tolerance $10^{-6}$. In the self-convergence test and for comparisons to the coarse BIE solution, 20 different problems are solved for each $\delta$, with randomly sampled no-slip boundary conditions.
Figure 14: Comparison to results from a BIE formulation for the tetrahedron test geometry of Figure 13(a); forces and torques determined with the image enhanced grid are compared to different sets of reference data. For small $\delta$, the BIE-reference is inaccurate due to the under-resolved peak in the layer density. This causes increasing errors for decreasing $\delta$ in comparisons of coarse MFS results both to solutions determined with coarse and fine BIE-grids (see details in main text), but with different onsets of the increase. For larger $\delta$, however, the error level in the forces and torques from the self-convergence test is similar in magnitude to the error in the comparison to a solution computed with BIE. For BIE, the QBX-tolerance is set to $10^{-6}$, which together with the GMRES tolerance restrict how small errors that can be detected for larger $\delta$. Red and blue represent the same type of comparisons, but using different reference solutions: for each separation distance $\delta$, the experiment is repeated 20 times with randomly generated boundary data, and the min, max and mean of $\epsilon_{\text{FT}}$ are visualized. Test problems generated with a unit force or torque are labeled ’bc A’ and ’bc B’. The self-convergence data is the same as presented in Figure 13(b).
Example 6 (Adaptive discretizations for randomized clusters).
We study random clusters of the type in Figure 1 consisting of 20 particles, with each sphere $\delta=10^{-3}$ away from at least one neighbor. In such a cluster, some spheres will have many close interactions, meaning that the adaptive number of image and collocation points will vary considerably between different particles. The spheres have randomly sampled rigid body velocities, with the resulting surface velocity exemplified in Figure 1(a) and relative residual over the sphere surfaces in Figure 1(b). Statistics for the maximum relative residuals, $\epsilon_{\text{res}}^{\max}$, for 50 realizations of the experiment, are shown in Figure 15(a). The same figure reports statistics for the number of close contacts for each realization. The adaptively determined number of image and collocation points for each sphere in each run are displayed in Figures 15(b) and 15(c), where the short bars for the largest number of points correspond to the few particles that have the largest number of near contacts in the center of the cluster. Note, however, that the maximum number of collocation points never exceeds the 5400 vector unknowns in the fine-grid BIE in Example 5 (being non-adaptive), and the solution with MFS is also much faster to compute. A majority of the particles have only one close neighbor, resulting in moderate numbers of unknowns, which is reflected in the tall bars in Figures 15(b) and 15(c).
(a) Statistics for the total number of close contacts (left) and the max relative residual (right).
(b) Statistics for the number of image points per sphere.
(c) Statistics for the number of collocation points per sphere.
Figure 15: By using an adaptive number of image points per closely interacting neighbor, as in (39), lubrication forces between $P=20$ spheres in a cluster with minimum separation $\delta=10^{-3}$ are resolved. Each sphere has a randomly sampled translational and angular velocity. The experiment is repeated 50 times. In panel (a), whiskers display min/max, box edges display the 10th and 90th percentile of results from the 50 runs, and the box center line displays the median.
5.3 GMRES iteration count
We have demonstrated controllable accuracy with our scheme on challenging problems with strong lubrication forces. In terms of efficiency, the computational grid can be on average much coarser than what is needed in a uniformly-discretized BIE method to resolve the interactions. It however remains to discuss the cost of iteratively solving the preconditioned system for small $\delta$. Here, we show for two of the example geometries from the previous section that the GMRES iteration count grows with decreasing particle distances, and that this phenomenon is common to both MFS and BIE methods.
Example 7 (continues=ex_BIE).
For the tetrahedron of spheres pair-wise separated by $\delta$, the GMRES iteration count as a function of $\delta$ is reported in Figure 16(a). The iteration count appears to grow like ${\mathcal{O}}(\delta^{-1/2})$ for all the tested boundary conditions (random velocity data, or, based on the reference solution, with either random forces and torques, or with a unit force/torque on each particle). The iteration counts required to provide the BIE reference with the fine grid are similar to those needed to compute the MFS solution with the same boundary condition. Note, however, that the numbers of BIE unknowns are larger – $5400\cdot 3$ per particle vs. $3M^{(i)}=3\cdot\left(801+12\cdot n_{c}\cdot 3\cdot 20\right)=3\cdot 2960$ with MFS, following (40). The displayed BIE GMRES iterations are for the inverse resistance problem (mobility), which we have found converges faster.
(a) GMRES iteration count for the data in Figure 14.
(b) GMRES iteration count for the data in Figure 15.
Figure 16: GMRES iteration count for the data in Examples 5 (panel (a)) and 6 (panel (b)), with the MFS GMRES tolerance set to $10^{-6}$. In panel (a) for the tetrahedron geometry, convergence is reported for different boundary data and compared to the convergence achieved with the BIE-formulation with QBX. The same trend of an increasing iteration count for decreasing $\delta$ is noted for both methods. The shaded data refers to max and min iteration counts corresponding to boundary conditions with randomly sampled rigid body velocities, 20 different sets for the fine grid and 40 sets for the coarse grid. Test problems generated with a unit force or torque are labeled ’bc A’ and ’bc B’ and converge slightly faster than problems with a boundary condition based on a randomly sampled velocity. Panel (b) displays iteration data for the clusters of 20 spheres with $\delta=10^{-3}$ and adaptively set image points, where whiskers display min/max, box edges the 10th and 90th percentile, and the box center lines display the median.
Example 8 (continues=adaptive).
Statistics for the GMRES count required for the randomized clusters of 20 spheres with adaptively set image points is reported in Figure 16(b).
To conclude, the GMRES count (for either MFS or BIE systems) becomes quite large for small $\delta$, and is consistent with ${\mathcal{O}}(\delta^{-1/2})$. Suitable preconditioning strategies are left to future work.
6 Conclusions and outlook
We consider exterior Stokes boundary value problems with no-slip boundary conditions applied to a collection of identical spherical particles immersed in a viscous fluid. The work presents a technique based on the method of fundamental solutions (MFS), with the flow in the basic setup represented as a linear combination of Stokeslets placed at proxy-surfaces in every particle interior. For problems with strong lubrication forces, this is augmented by a combination of Stokeslets, rotlets and potential dipoles that discretize each pair-wise “image line”, providing controllable residuals for all separations down to $\delta=10^{-3}$ for spheres of unit radii. We present a recipe, as a function of such a $\delta$, for the number of image points and accompanying “caps” of collocation nodes, to reach a target accuracy of $10^{-3}$ uniformly in the surface velocities. We show that, remarkably, such line-image approximations developed for pairwise interactions also perform equally well for packed clusters with many close spheres.
Interactions for close-to-touching particles moving relative to each other are challenging to compute with established methods, such as a boundary integral (BIE) method with uniform Nyström quadrature, because a sharply peaked force density has to be resolved. With the MFS technique, we demonstrate numerically that lubrication forces can be accurately resolved using much coarser particle representations (less unknowns) than a uniform BIE method at the same accuracy. It would be interesting to compare against spatially-adaptive BIE quadratures, which are less established. In the present work, by enforcing boundary conditions by collocation at the particle surfaces in the least-squares sense, the technical singular quadratures needed for BIE are avoided. In the special case without lubrication forces due to relative particle motions, with a backward-stable one-body preconditioning strategy, plain proxy sources are sufficient to reach the target accuracy. The sum of Stokeslets can then be rapidly evaluated using the fast multipole method, yielding close to linear scaling in the number of particles, which we demonstrate for systems with up to 4.8 million degrees of freedom.
A natural extension to the work reported here is to the mobility problem. We report on a well-conditioned MFS formulation for the mobility problem separately in [75]. In that work, we also extend to non-spherical smooth particles in the setting where proxy surfaces suffice to resolve interactions. It is a topic of future research whether image representations could be useful for non-spherical particles. In Section 5.3, we note that the number of GMRES iterations grows with decreasing particle gaps. Future work includes investigating suitable preconditioning or the use of a fast direct solver. As for the size of the systems considered in this work, all tests fit on a standard workstation. Larger systems could be studied via a distributed memory implementation and FMM, as in [30].
Acknowledgements
Broms and Tornberg acknowledge the support from the Swedish Research Council: grant no. 2019-05206 and the research environment grant INTERFACE (biomaterials), no. 2016-06119. The Flatiron Institute is a division of the Simons Foundation. Broms appreciates the research visit at CCM Flatiron, which served as a catalyst for the initial ideas in this work. We would also like to thank the referees for their thorough reading of the manuscript and for their valuable comments which have helped us to improve the text.
Appendix A Singular values
As for the numerical examples in Section 5.1, the coarse MFS grid with $N=686$ and $R_{p}=0.63$ is here used for numerical experiments. Further, for a sphere close-to-touching with one other sphere such that image points are used, $M=1.2\cdot N+2\cdot 6\cdot 3n_{\text{im}}$.
A.1 Truncation of singular values with image enhancement
It is difficult to get the approximation power of the image points as large as possible for a general choice of no-slip boundary conditions, while keeping the ill-conditioning under control. Truncation of singular values of the self-interaction blocks in the target-from-source matrix is hence necessary and the effect of the truncation level, $\epsilon_{\text{trunc}}$, relative to the largest singular value, is here investigated.
As discussed in Section 2.1, clustering of source points worsens the conditioning of the self-interaction blocks in the target-from-source matrix. Two clustering properties of the image points defined by (26)-(27) impact this. As $\delta$ decreases, image points approach the particle surface ($R_{\text{acc}}(\delta)\to 1$), making the fundamental solutions near-singular when evaluated near the closest approach points, while for larger $\delta$, image points are further from the surface, reducing near-singularity issues. Additionally, with Chebyshev sampling and equal $n_{\text{im}}$, image points are more clustered for larger $\delta$, increasing ill-conditioning.
For a fixed no-slip boundary condition on two spheres $\delta$ apart, the maximum relative residual, $\epsilon_{\text{res}}^{\max}$, as defined by (30), is computed as a function of the number of image points, $n_{\text{im}}$, for a few different choices of $\epsilon_{\text{trunc}}$. The result is visualized in Figure 17 for $\delta=10^{-2}$ and $\delta=10^{-3}$; high truncation generally results in larger residuals but smaller coefficient magnitudes. Given experiments of this type, we pick $\epsilon_{\text{trunc}}=5\cdot 10^{-12}$ for the numerical examples presented in the results section of the paper, Section 5.
(a) $\delta=10^{-2}$, max relative residual $\epsilon_{\text{res}}^{\max}$.
(b) $\delta=10^{-2}$, max coefficient magnitude $\|\boldsymbol{\lambda}\|_{\infty}$.
(c) $\delta=10^{-3}$, max relative residual $\epsilon_{\text{res}}^{\max}$.
(d) $\delta=10^{-3}$, max coefficient magnitude $\|\boldsymbol{\lambda}\|_{\infty}$.
Figure 17: Effect of truncation of singular values in the self-interaction blocks of the target-from-source matrix, investigated for two spheres separated by $\delta$. Notably, $\epsilon_{\text{res}}^{\max}$ decreases and $\|\boldsymbol{\lambda}\|_{\infty}$ stays bounded with the number of image points per particle $n_{\text{im}}$, only if the smallest singular values are truncated. However, high truncation of singular values come at a cost, as information then is lost. A reasonable level of truncation is chosen to be $\epsilon_{\text{trunc}}=5\cdot 10^{-12}$.
A.2 Decay of the singular value spectrum
Fixing $N=686$ proxy-sphere points and no added image points, Figure 18(a) plots the singular value spectrum for the target-from-source (self-interaction) matrix denoted $\boldsymbol{B}$ in Section 2.2. The $j$th singular value is well-predicted as $\mathcal{O}(R_{p}^{c\sqrt{j}})$; see Remark 2. Hence, the ill-conditioning is more severe, the smaller the radius of the proxy-surface, $R_{p}$. Then adding images, singular values for one particle in a pair separated by $\delta$ are displayed in Figures 18(c) and 18(b) without and with left-preconditioning, as described in Remark 5 of Section 5. The more closely interacting the particles are, the more ill-conditioned is the self-interaction matrix. Left-preconditioning does not remove the numerical nullspace, but does decrease the largest few singular values by 1–2 orders of magnitude.
(a) Decay of singular values for a single particle without image points.
(b) With image points: Example without preconditioning from the left.
(c) With image points: Example with preconditioning from the left.
Figure 18: Decay of singular values for one particle with a single close neighbor at a distance $\delta$. In panel (a), without images, the radius of the proxy-surface, $R_{p}$, has an impact on the magnitude of the singular values; see also the discussion in Section 2.1. With image points in panels (b) and (c), the magnitude of the largest singular values are increased, the smaller $\delta$ is. These singular values are in panel (c) decreased by preconditioning from the left.